A farsighted person cannot see objects placed closer to 50 cm. Find the power of the lens needed to see the objects at 20 cm.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. For the far sighted person, u = – 20 cm, v = – 50 cm from lens formula 1/v – 1/u = 1/f 1/f = 1/-50 – 1/-20 = 1/20 – 1/50 = 3/100 ⇒ f = 100/3 cm = 1/3 m So, power of the lens = 1/f = 3 Diopter